# Prior-based Noisy Text Data Filtering: Fast and Strong Alternative For Perplexity

### v2

@@ 1 Introduction
To address this need, various data selection methods have been proposed. Early approaches relied on heuristic rules [33, 5], but more recent trends have shifted toward model-based techniques [44, 23]. These methods typically involve training a reference model on a target dataset and using it to identify desirable data. The model may perform binary classification [43] or compute similarity with the reference dataset [44]. Among these, using the perplexity (PPL) score from a reference model as a criterion of filtering is currently known to offer the best performance while maintaining a relatively simple implementation [4]. We provide a more detailed review of related work in §A.
However, PPL-based approaches come with the following inherent limitations. (1) Time cost: These methods require training a reference model, followed by inference of PPL over the whole corpus. Given that web-scale data can easily exceed trillions of documents and continues to grow, performing inference over the entire corpus becomes prohibitively expensive. (2) Reliability: LLMs often fail to accurately assess samples from distributions that is not seen while training, such as noisy data. As a result, generative perplexity may sometimes assign high scores to noisy or low-quality text [14, 42]. This issue might become more pronounced when using smaller models to reduce inference costs, further undermining reliability.
@@
Interestingly, this method is inspired by traditional techniques used in deciphering ancient languages. The 8th-century linguist Al-Kindi first proposed that, in order to decipher an encrypted language, analyzing the frequency of its words provides a clue [2]. If some word appears with the highest frequency across multiple documents, it is likely to correspond to a function word, such as "is" or "a" in English. This indicates that term-frequency itself is a one-dimensional representation for the role of a word: high frequency maps to function words while relatively low frequency maps to content words (e.g., “US”, “president”). Combining with another linguistic observation that well-formed sentences within a language tend to exhibit a consistent level of lexical density (i.e., ratio between function and content words) [16], we can determine outlier document simply by computing the mean and variance of its term frequencies: which we term prior-based data filter.
@@ 2.1 PPL-based approach and estimation of prior
The PPL-based filtering method is known as the most effective approach for filtering noise data from web text corpus for pretraining LLMs [4, 23]. For the filtering, first, a small reference model $\theta$ (an autoregressive transformer architecture of 137M parameters) is trained on the corpus $D$. The model then computes the PPL for each data point $\mathtt{d}=(x_{1},x_{2},\dots,x_{N})$, where $x_{i}$ is the token at the $i^{th}$ position of a document, and $\mathtt{d}\in D$. Then, $\mathtt{d}$ with PPL values farthest from the median are discarded. Here, the PPL is defined as follows:
@@ 2.2 Frequency analysis in linguistics
(1) Term frequency is a 1-dimensional representation of a word’s role: The 8th-century linguist Al-Kindi first proposed an idea that is still widely used today [2]: to decipher ancient or encrypted languages, analyzing the frequency of its words gives a clue. If some word appears with the highest frequency across multiple documents, it is likely to correspond to a function word (e.g., "is" or "a" in English) that serves grammatical roles. In contrast, content words which carry semantic meaning (e.g., “US”, “president”) tend to appear with relatively lower frequency. Therefore, frequency itself can serve as a basis for distinguishing between function words and content words. In other words, term frequency (i.e., prior) can be seen as a one-dimensional representation of a word’s functional role. We analyze that this property partially stems from the next property.
(2) Well-formed sentences typically exhibit a consistent range of lexical density: As lexical density is defined as the proportion of content words against function words, it is known that well-formed sentences in a language typically maintain a certain range of lexical density [16]. From this, we can infer that broken and ill-formed sentences will deviate significantly from this range to be outliers.
@@ 3.1 Analysis on the token prior
We first analyze the token-level term frequencies by calculating the token priors (with formulation in §2.1) on the Dolma dataset [39]. As we sort them by the logarithm (Figure 1), we can observe that the token priors distinctly fall into three clusters based on their height and slope, supporting the thesis that the token prior serves as a 1-dimensional representation for the token’s role.
@@ 3.4.1 Prior-based filter approximates PPL-based filter
$\footnotesize\log\text{PPL}({\mathtt{d}})\propto\underbrace{\sum^{N}_{i}\log p_{\theta}(x_{<i}|x_{i})}_{\pi_{likelihood}}+\underbrace{\sum_{i}^{N}\log p_{\theta}(x_{i})}_{\pi_{prior}}$ (4)
@@
Prior can be even better metric than PPL While $\sigma_{d}$ captures an approximation of the likelihood term, it is significantly more saturated than the actual likelihood, which can be considered a limitation. However, conversely, the inherent instability of the $\pi_{likelihood}$ (described as follows) poses a limitation for the PPL-based approach. (1) When the model is small, it struggles to accurately learn the likelihood [41]. (2) The model does not learn how to estimate likelihood for data from previously unseen distributions (mostly noisy data), which is not a problem for estimating only the prior. For this reason, previous studies have also reported that PPL often mistake repetitive or pattern-based noise as valid text [14]. Empirically, the model trained with the prior-based filter shows better downstream performance than the one trained with the PPL-based filter (§4).
@@
Observation on filtered samples This characteristic of PPL is also observed in outlier samples. We investigate the most extreme outlier samples from each metric (PPL, $\mu_{\mathtt{d}}$, $\sigma_{\mathtt{d}}$), excluding their overlaps (Figure 3). As described in §3.3, outliers of $\mu_{\mathtt{d}}$ tend to be filled with extremely low or high prior tokens (Figure 3b), while those of $\sigma_{\mathtt{d}}$ often consist of content words but lack function words or valid sentence structure (Figure 3c). In outliers of PPL (Figure 3a), content and function words appear to be well-balanced, giving the surface impression of well-formed sentences, but upon closer reading, many of them turn out to be semantically meaningless. This may reveal both a strength and a weakness of the PPL metric: it effectively captures subtle irregularities within well-formed documents, but may fail to detect noise arising from entirely out-of-distribution samples.
@@
The results show a strong correlation: When filtering by the $e=0.10$, nearly 50% of $F_{\mu}$ and $F_{\text{ppl}}$ overlap. We also find $F_{\mu}$ aligns more closely with $F_{\text{ppl}}$ than $F_{\sigma}$. We provide additional evidence in §B.3
@@ 3.4.2 $\mu_{\mathtt{d}}$ reflects learnability of minor language in multi-lingual setting
This characteristic offers a major advantage over methods that require manually specifying a reference dataset (e.g., DSIR [44]). In DSIR, a human must decide whether to select English or Chinese data and then provide a suitable reference dataset accordingly. In contrast, the prior-based filter automatically determines whether a language should be filtered out based on its learnability.
@@ 4 Experiment on downstream task
In this section, we evaluate the downstream task performance of models pretrained with different data filtering methods. Most training settings and hyperparameters follow those of “Perplexed by perplexity: Perplexity-based data pruning with small reference model [4]”. We first conduct experiments on a natural language (specified to English) web corpus, Dolma [39]. This allows us to assess the effects on general language capabilities of a model (e.g., knowledge, language understanding, and symbolic understanding). To demonstrate that our prior-based method is applicable even to symbolic languages such as code and math, we also perform experiments on the Pile-github [^2] https://www.kaggle.com/datasets/dschettler8845/the-pile-github-files-part-01 dataset.
@@ 4.1 Experiment on natural language corpus and general ability
Following [4], we mainly use Dolma [39] as a pretraining corpus for a testbed of filtering methods. Dolma is a large-scale, diverse web-text corpus, designed for training and evaluating LLMs. It contains noisy web data sources that support general language use ability, such as world knowledge, commonsense reasoning, and symbolic problem solving. This corpus is composed of multiple web-scale datasets, including Common Crawl, Reddit, Wikipedia, and Wikibooks [^3] https://commoncrawl.org/, https://www.reddit.com/, https://dumps.wikimedia.org/, The Stack [19], C4 [33], PeS2o [38], Project Gutenberg [3] (see Table 1). Among these, Common Crawl accounts for the major portion (74.5%) of the corpus. This makes it a particularly suitable environment for evaluating filtering methods, as it contains a high proportion of noisy web content that must be thoroughly filtered, while a small but valuable subset (e.g., books and educational data) must be preserved. For testing under resource constraints, we select v1.6—a smaller subset with 6.3B tokens. We divide this into blocks ($\mathtt{d}$}) of 512 tokens, and select a subset of $N$ tokens for pretraining.
@@
Baseline setup When selecting a subset from Dolma, we follow the procedure defined by each method: (1) no-filter: Randomly selects $N$ without applying any filtering method. (2) PPL-based: Following the approach of [4] and §2.1, we first train a reference model (137M) on the random 3B tokens subset of dataset. We then compute the PPL score for each sample in the dataset. To obtain a final subset of size $N$, we discard samples with the highest and lowest PPL scores. (3) DSIR: Adopting the well-known method DSIR [44], we estimate n-gram frequency from the reference dataset (we choose Bookcorpus and Wiki-en) and compute importance weights. DSIR is proposed as an advancement of FastText, a classifier-based method using manual data curation. (4) prior-based (ours): As described in §3.2, we first estimate token priors using a 10% subset of the full corpus. Based on these priors, we compute $\mu_{d}$ and $\sigma_{d}$ (${\mathtt{d}}\in D$). We then discard samples with the highest $\delta_{\mu}$ and $\delta_{\sigma}$ values in the constraint of $|F_{\mu}|=|F_{\sigma}|$, until the volume of final subset $|F_{\mu}\cup F_{\sigma}|$ reaches $N$. For all baselines, N is set to 50% (3B), which is observed in [4] to yield the best performance.
@@
We use the GPT-2 architecture for pretraining, with large (1.5B) and small (137M) size models, using 8 GPUs (RTX A5000). Following [4], we set a max token length of 512, a global batch size 256, and a learning rate 2e-4, and train for 40K global steps (about 6B token duration). According to [4], the relative performance trends observed at 40K steps are maintained in later training steps.
As our study requires a significant amount of resources for pre-training across multiple baselines, we carefully adopted settings and observations from recent representative works to ensure experimental rigor under limited resources. For example, [44, 4, 22] empirically shows that DSIR and PPL-based filtering mostly perform better than other rule-based (e.g., filtering based on character length, inclusion of certain words, and the ratio of English content) and model-based (e.g., FastText, memorization score, classifier on LM embeddings) methods. We therefore regard DSIR and PPL-based filtering as representing the most advanced methods to date. Further explanation is in §A and §C.2.
@@
Benchmark and evaluation setup The types and settings of downstream tasks follow those used in the [4], based on the MosaicML evaluation gauntlet [25]. Gauntlet includes tasks designed to assess five core capabilities: world knowledge, common sense reasoning, language understanding, symbolic problem solving, and reading comprehension. We normalize the accuracy of the individual task as $a_{n}=\frac{a_{m}-a_{r}}{1-a_{r}}$, where $a_{m}$ is the accuracy of the model and $a_{r}$ is the expected accuracy of random guessing. We report the average normalized accuracy for each task, task category, and the average across all categories. Since some tasks are not proper for 1.5B models, we exclude benchmarks with average $a_{n}$ of baselines under 0.001. This results in a total of 20 benchmarks (details in §D)
@@
Table 2: Performance and time cost (for filtering) of the baselines pre-trained on Dolma across 20 benchmarks. The average normalized accuracy is the average of all categories.
Time
Average
normalized
accuracy
World
knowledge
Commonsense
reasoning
Language
understanding
Symbolic
problem
solving
Reading
compre-
hension
Large (1.5B) model
   no-filter - 5.78 5.52 0.44 6.14 13.22 3.59
   DSIR 4 hours 7.56 7.03 6.84 7.31 12.67 3.97
   PPL-based 216216 GPU hours 8.22 9.98 11.91 7.34 7.91 3.96
   Prior-based (ours) 0.25 hours 9.20 9.53 11.27 10.31 11.13 3.79
Small (137M) model
   no-filter - 4.96 4.96 1.81 1.47 12.83 3.70
   DSIR 4 hours 5.60 5.68 4.93 1.97 11.60 3.80
   PPL-based 216216 GPU hours 5.26 5.47 6.53 2.90 7.84 3.58
@@
Results As described in Table 2, the results show that the model trained with prior-based filtering achieves the highest average performance, with extremely small time cost. Key observations are as follows: (1) DSIR outperforms no-filter, and PPL-based outperforms DSIR, which aligns with findings from previous research [4, 44]. (2) Prior-based filter approximates PPL-based filter in principle, but yields better downstream performance. We analyze that this is because PPL score depends on the model’s likelihood, which can be unstable. On the other hand, the prior is based on simple word frequencies, so it gives a more stable and reliable signal. (3) Though the prior-based model outperforms the PPL-based model in downstream performance, the prior-based filtering requires significantly less processing time. PPL-based filtering takes 216 GPU hours to select a 3B token subset ($20\times 8$ GPU hours of training the reference model, $7\times 8$ GPU hours of PPL inference), while prior-based filtering takes only 15 minutes (6 minutes of assessing token prior, 6 minutes of calculating $\mu_{\mathtt{d}}$ and $\sigma_{\mathtt{d}}$ in DD)—under 0.1% of the time spent for PPL. This demonstrates the superior scalability and efficiency of our approach.
@@
We provide further analyses and ablations for key issues in §B, such as preserving minority data, and sensitivity to factors (e.g., tokenizer, block size).

### v4

@@ 1 Introduction
To address this need, various data selection methods have been proposed. Early approaches relied on heuristic rules (Raffel et al., 2023; Bane et al., 2022), but more recent trends have shifted toward model-based techniques (Xie et al., 2023b; Marion et al., 2023). These methods typically involve training a reference model on a target dataset and using it to identify desirable data. The model may perform binary classification (Xie et al., 2023a) or compute similarity with the reference dataset (Xie et al., 2023b). Among these, using the perplexity (PPL) score from a reference model as a criterion of filtering is currently known to offer the best performance while maintaining a relatively simple implementation (Ankner et al., 2024). We provide a more detailed review of related work in §A.
However, PPL-based approaches come with the following inherent limitations. (1) Time cost: These methods require training a reference model, followed by inference of PPL over the whole corpus. Given that web-scale data can easily exceed trillions of documents and continues to grow, performing inference over the entire corpus becomes prohibitively expensive. (2) Reliability: LLMs often fail to accurately assess samples from distributions that is not seen while training, such as noisy data. As a result, generative perplexity may sometimes assign high scores to noisy or low-quality text (Holtzman et al., 2020; Welleck et al., 2019). This issue might become more pronounced when using smaller models to reduce inference costs, further undermining reliability.
@@
Interestingly, this method is inspired by traditional techniques used in deciphering ancient languages. The 8th-century linguist Al-Kindi first proposed that, in order to decipher an encrypted language, analyzing the frequency of its words provides a clue (Al-Kadit, 1992). If some word appears with the highest frequency across multiple documents, it is likely to correspond to a function word, such as "is" or "a" in English. This indicates that term-frequency itself is a one-dimensional representation for the role of a word: high frequency maps to function words while relatively low frequency maps to content words (e.g., “US”, “president”). Combining with another linguistic observation that well-formed sentences within a language tend to exhibit a consistent level of lexical density (i.e., ratio between function and content words) (Johansson, 2008), we can determine outlier document simply by computing the mean and variance of its term frequencies: which we term prior-based data filter.
@@ 2.1 PPL-based approach and estimation of prior
The PPL-based filtering method is known as the most effective approach for filtering noise data from web text corpus for pretraining LLMs (Ankner et al., 2024; Marion et al., 2023). For the filtering, first, a small reference model $\theta$ (an autoregressive transformer architecture of 137M parameters) is trained on the corpus $D$. The model then computes the PPL for each data point $\mathtt{d}=(x_{1},x_{2},\dots,x_{N})$, where $x_{i}$ is the token at the $i^{th}$ position of a document, and $\mathtt{d}\in D$. Then, $\mathtt{d}$ with PPL values farthest from the median are discarded. Here, the PPL is defined as follows:
@@ 2.2 Frequency analysis in linguistics
(1) Term frequency is a 1-dimensional representation of a word’s role: The 8th-century linguist Al-Kindi first proposed an idea that is still widely used today (Al-Kadit, 1992): to decipher ancient or encrypted languages, analyzing the frequency of its words gives a clue. If some word appears with the highest frequency across multiple documents, it is likely to correspond to a function word (e.g., "is" or "a" in English) that serves grammatical roles. In contrast, content words which carry semantic meaning (e.g., “US”, “president”) tend to appear with relatively lower frequency. Therefore, frequency itself can serve as a basis for distinguishing between function words and content words. In other words, term frequency (i.e., prior) can be seen as a one-dimensional representation of a word’s functional role. We analyze that this property partially stems from the next property.
(2) Well-formed sentences typically exhibit a consistent range of lexical density: As lexical density is defined as the proportion of content words against function words, it is known that well-formed sentences in a language typically maintain a certain range of lexical density (Johansson, 2008). From this, we can infer that broken and ill-formed sentences will deviate significantly from this range to be outliers.
@@ 3 Prior-based data filtering
In §C, we further provide a theoretical and empirical explanation that the prior-based filter generalizes across diverse types of languages and tokenizers.
@@ 3.1 Analysis on the token prior
We first analyze the token-level term frequencies by calculating the token priors (with formulation in §2.1) on the Dolma dataset (Soldaini et al., 2024). As we sort them by the logarithm (Figure 1), we can observe that the token priors distinctly fall into three clusters based on their height and slope, supporting the thesis that the token prior serves as a 1-dimensional representation for the token’s role.
@@ 3.2 Formulation of prior-based data filtering
In §B, we provide a theoretical derivation of the noise-detection behavior induced by $\mu_{\mathtt{d}}$ and $\sigma_{\mathtt{d}}$, and show that it approximates that of a PPL-based filter.
@@ 3.4.1 Prior-based filter approximates PPL-based filter
$\log\text{PPL}({\mathtt{d}})\propto\underbrace{\sum^{N}_{i}\log p_{\theta}(x_{<i}|x_{i})}_{\pi_{likelihood}}+\underbrace{\sum_{i}^{N}\log p_{\theta}(x_{i})}_{\pi_{prior}}$ (4)
@@
Prior can be even better metric than PPL While $\sigma_{d}$ captures an approximation of the likelihood term, it is significantly more saturated than the actual likelihood, which can be considered a limitation. However, conversely, the inherent instability of the $\pi_{likelihood}$ (described as follows) poses a limitation for the PPL-based approach. (1) When the model is small, it struggles to accurately learn the likelihood (Wei et al., 2022). (2) The model does not learn how to estimate likelihood for data from previously unseen distributions (mostly noisy data), which is not a problem for estimating only the prior. For this reason, previous studies have also reported that PPL often mistake repetitive or pattern-based noise as valid text (Holtzman et al., 2020). Empirically, the model trained with the prior-based filter shows better downstream performance than the one trained with the PPL-based filter (§4).
@@
In §B, we provide a theoretical explanation of how the prior-based filter approximates the noise-detection behavior of the PPL-based filter, as well as the boundary conditions where a discrepancy between the two emerges. We also explain why median-based thresholding is robust.
Observation on filtered samples
This characteristic of PPL is also observed in outlier samples. We investigate the most extreme outlier samples from each metric (PPL, $\mu_{\mathtt{d}}$, $\sigma_{\mathtt{d}}$), excluding their overlaps (Figure 3). As described in §3.3, outliers of $\mu_{\mathtt{d}}$ tend to be filled with extremely low or high prior tokens (Figure 3b), while those of $\sigma_{\mathtt{d}}$ often consist of content words but lack function words or valid sentence structure (Figure 3c). In outliers of PPL (Figure 3a), content and function words appear to be well-balanced, giving the surface impression of well-formed sentences, but upon closer reading, many of them turn out to be semantically meaningless. This may reveal both a strength and a weakness of the PPL metric: it effectively captures subtle irregularities within well-formed documents, but may fail to detect noise arising from entirely out-of-distribution samples.
@@
The results show a strong correlation: When filtering by the $e=0.10$, nearly 50% of $F_{\mu}$ and $F_{\text{ppl}}$ overlap. We also find $F_{\mu}$ aligns more closely with $F_{\text{ppl}}$ than $F_{\sigma}$. We provide additional evidence in §E.2
@@ 3.4.2 $\mu_{\mathtt{d}}$ reflects learnability of minor language in multi-lingual setting
This characteristic offers a major advantage over methods that require manually specifying a reference dataset (e.g., DSIR (Xie et al., 2023b)). In DSIR, a human must decide whether to select English or Chinese data and then provide a suitable reference dataset accordingly. In contrast, the prior-based filter automatically determines whether a language should be filtered out based on its learnability.
@@ 4 Experiment on downstream task
In this section, we evaluate the downstream task performance of models pretrained with different data filtering methods. Most training settings and hyperparameters follow those of “Perplexed by perplexity: Perplexity-based data pruning with small reference model (Ankner et al., 2024)”. We first conduct experiments on a natural language (specified to English) web corpus, Dolma (Soldaini et al., 2024). This allows us to assess the effects on general language capabilities of a model (e.g., knowledge, language understanding, and symbolic understanding). To demonstrate that our prior-based method is applicable even to symbolic languages such as code and math, we also perform experiments on the Pile-github [^3] https://www.kaggle.com/datasets/dschettler8845/the-pile-github-files-part-01 dataset.
@@ 4.1 Experiment on natural language corpus and general ability
Following Ankner et al. (2024), we mainly use Dolma (Soldaini et al., 2024) as a pretraining corpus for a testbed of filtering methods. Dolma is a large-scale, diverse web-text corpus, designed for training and evaluating LLMs. It contains noisy web data sources that support general language use ability, such as world knowledge, commonsense reasoning, and symbolic problem solving. This corpus is composed of multiple web-scale datasets, including Common Crawl, Reddit, Wikipedia, and Wikibooks [^4] https://commoncrawl.org/, https://www.reddit.com/, https://dumps.wikimedia.org/, The Stack (Kocetkov et al., 2022), C4 (Raffel et al., 2023), PeS2o (Soldaini and Lo, 2023), Project Gutenberg (Angelescu, 2013) (see Table 1). Among these, Common Crawl accounts for the major portion (74.5%) of the corpus. This makes it a particularly suitable environment for evaluating filtering methods, as it contains a high proportion of noisy web content that must be thoroughly filtered, while a small but valuable subset (e.g., books and educational data) must be preserved. For testing under resource constraints, we select v1.6—a smaller subset with 6.3B tokens. We divide this into blocks ($\mathtt{d}$}) of 512 tokens, and select a subset of $N$ tokens for pretraining.
@@
Baseline setup When selecting a subset from Dolma, we follow the procedure defined by each method: (1) no-filter: Randomly selects $N$ without applying any filtering method. (2) PPL-based: Following the approach of Ankner et al. (2024) and §2.1, we first train a reference model (137M) on the random 3B tokens subset of dataset. We then compute the PPL score for each sample in the dataset. To obtain a final subset of size $N$, we discard samples with the highest and lowest PPL scores. (3) DSIR: Adopting the well-known method DSIR (Xie et al., 2023b), we estimate n-gram frequency from the reference dataset (we choose Bookcorpus and Wiki-en) and compute importance weights. DSIR is proposed as an advancement of FastText, a classifier-based method using manual data curation. (4) prior-based (ours): As described in §3.2, we first estimate token priors using a 10% subset of the full corpus. Based on these priors, we compute $\mu_{d}$ and $\sigma_{d}$ (${\mathtt{d}}\in D$). We then discard samples with the highest $\delta_{\mu}$ and $\delta_{\sigma}$ values in the constraint of $|F_{\mu}|=|F_{\sigma}|$, until the volume of final subset $|F_{\mu}\cup F_{\sigma}|$ reaches $N$. For all baselines, N is set to 50% (3B), which is observed in Ankner et al. (2024) to yield the best performance.
@@
We use the GPT-2 architecture for pretraining, with large (1.5B) and small (137M) size models, using 8 GPUs (RTX A5000). Following Ankner et al. (2024), we set a max token length of 512, a global batch size 256, and a learning rate 2e-4, and train for 40K global steps (about 6B token duration). According to Ankner et al. (2024), the relative performance trends observed at 40K steps are maintained in later training steps.
As our study requires a significant amount of resources for pre-training across multiple baselines, we carefully adopted settings and observations from recent representative works to ensure experimental rigor under limited resources. We explain how our baseline lineup comprehensively covers the major streams of previous works in §A.2. We also provide justification of our experimental scale (e.g., training token duration, model size) in §F.2.
@@
Benchmark and evaluation setup The types and settings of downstream tasks follow those used in the Ankner et al. (2024), based on the MosaicML evaluation gauntlet (MosaicML, 2023). Gauntlet includes tasks designed to assess five core capabilities: world knowledge, common sense reasoning, language understanding, symbolic problem solving, and reading comprehension. We normalize the accuracy of the individual task as $a_{n}=\frac{a_{m}-a_{r}}{1-a_{r}}$, where $a_{m}$ is the accuracy of the model and $a_{r}$ is the expected accuracy of random guessing. We report the average normalized accuracy for each task, task category, and the average across all categories. Since some tasks are not proper for 1.5B models, we exclude benchmarks with average $a_{n}$ of baselines under 0.001. This results in a total of 20 benchmarks (details in §G)
Results As described in Table 2, the results show that the model trained with prior-based filtering achieves the highest average performance, with extremely small time cost. Key observations are as follows: (1) DSIR outperforms no-filter, and PPL-based outperforms DSIR, which aligns with findings from previous research (Ankner et al., 2024; Xie et al., 2023b). (2) Prior-based filter approximates PPL-based filter in principle, but yields better downstream performance. We analyze that this is because PPL score depends on the model’s likelihood, which can be unstable. On the other hand, the prior is based on simple word frequencies, so it gives a more stable and reliable signal. (3) Though the prior-based model outperforms the PPL-based model in downstream performance, the prior-based filtering requires significantly less processing time. PPL-based filtering takes 216 GPU hours to select a 3B token subset ($20\times 8$ GPU hours of training the reference model, $7\times 8$ GPU hours of PPL inference), while prior-based filtering takes only 15 minutes (6 minutes of assessing token prior, 6 minutes of calculating $\mu_{\mathtt{d}}$ and $\sigma_{\mathtt{d}}$ in DD)—under 0.1% of the time spent for PPL. This demonstrates the superior scalability and efficiency of our approach.
@@
Table 2: Performance and time cost (for filtering) of the baselines pre-trained on Dolma across 20 benchmarks. The average normalized accuracy is the average of all categories. Avg: Avg normalized accuracy, W: World knowledge, C: Commonsense reasoning, L: Language understanding, S: Symbolic problem solving, R: Reading comprehension
Type Time Avg W C L S R
1.5B model
   vanilla rule-based - 5.78 5.52 0.44 6.14 13.22 3.59
   FastText classifier 3.6 hrs 7.09 6.71 6.11 6.89 11.93 3.82
   DSIR n-gram 4 hrs 7.56 7.03 6.84 7.31 12.67 3.97
   PPL-based model-based 216216 GPU hrs 8.22 9.98 11.91 7.34 7.91 3.96
   Prior-based (ours) 0.25 hrs 9.20 9.53 11.27 10.31 11.13 3.79
137M model
   vanilla rule-based - 4.96 4.96 1.81 1.47 12.83 3.70
   FastText classifier 3.6 hrs 5.39 5.12 4.29 1.74 12.31 3.49
   DSIR n-gram 4 hrs 5.60 5.68 4.93 1.97 11.60 3.80
   PPL-based model-based 216216 GPU hours 5.26 5.47 6.53 2.90 7.84 3.58
@@
Further analysis
As data filtering removes noisy data, which is often found in outliers, it inherently risks discarding valuable minority data (e.g., a highly technical paper with rare terminology, a corpus including minority languages) as a trade-off. In §D, we analyze the robustness of the prior-based filter with respect to this trade-off.
In §C, we further provide a theoretical and empirical explanation that the prior-based filter generalizes across diverse types of languages and tokenizers. We also provide analysis on the sensitivity to factors such as block size and threshold in §E.
4.1.1 Consistency in a large-scale setting
We conduct experiments on larger models and datasets. For the efficiency of massive experiments, we adopt the strongest baselines from previous experiments (PPL-based and prior-based). We evaluate 3B (Qwen2.5-3B) and 1.5B models with training up to 12B token durations ($2\times$ repetitions). The results consistently demonstrate the superiority of the prior-based filter (Figure 7).
(a) 3B model
(b) 1.5B model
Figure 7: Average normalized accuracy across 20 benchmarks.